# FairEM360: A Suite for Responsible Entity Matching

### v1

@@ Abstract.
Entity matching is one the earliest tasks that occur in the big data pipeline and is alarmingly exposed to unintentional biases that affect the quality of data. Identifying and mitigating the biases that exist in the data or are introduced by the matcher at this stage can contribute to promoting fairness in downstream tasks. This demonstration showcases FairEM360, a framework for 1) auditing the output of entity matchers across a wide range of fairness measures and paradigms, 2) providing potential explanations for the underlying reasons for unfairness, and 3) providing resolutions for the unfairness issues through an exploratory process with human-in-the-loop feedback, utilizing an ensemble of matchers. We aspire for FairEM360 to contribute to the prioritization of fairness as a key consideration in the evaluation of EM pipelines.
$*{}^{*}$uthors contributed equally to this work.
@@ 1. Introduction
FairEM360 incorporates a comprehensive set of group fairness definitions tailored for single and pairwise fairness metrics specifically designed for EM audits (Shahbazi et al., 2023a). FairEM360 also features a ready-to-use ensemble of 10 matchers for performing matching tasks. Furthermore, FairEM360 offers a diverse range of subgroup-based, measure-based, and group-representation based explanations that shed light on the reasoning behind matcher’s unfair behavior. Finally, through an exploratory process with human-in-the-loop feedback that navigates the combinatorial space of matchers, and fairness measures, FairEM360 is directed toward satisfying the fairness constraints required by the user while satisfying a minimum matching performance for various groups. To the best of our knowledge, FairEM360 stands out as the first system for responsible EM.
@@ 2.1. Data Layer
Group Extraction: The first step in auditing an entity matcher for fairness is identifying meaningful groups/subgroups (e.g., white-male, black-female, ect.) from the sensitive attributes (e.g., race, sex), based on which the matcher should be audited. Sensitive attributes are attributes for which a matcher is likely to exhibit bias. Depending on the type, cardinality, and number of sensitive attributes, FairEM360 navigates the space of all possible (sub)groups. It unifies the group representations using one-hot encoding, which enables representing the subgroups of individual entities and pairs of entities across various settings such as binary, non-binary, and setwise sensitive attributes. Entity encodings are the output of the data layer and will be passed as an input to the logic layer, where the fairness of a matcher is investigated for a group.
Preprocessing: If the user opts to utilize the integrated matchers in their evaluation, the data layer will be responsible for preprocessing the input datasets to ensure their compatibility with the matchers, as well as splitting the data into test, train, and validation sets. The datasets are passed to the logic layer to be trained.
@@ 2.2. Logic Layer
The actual evaluation of the output from an EM task occurs in the logic layer. The input to the logic layer is a workload of tuples each consisting of left and right entity groups (extracted in the data layer), the matcher prediction, and the matching ground truth. A workload is a test set of tuples for evaluating an entity matcher. Given the pairwise nature of EM tasks, single fairness or pairwise fairness paradigms can be used to audit matchers for fairness. In single fairness, the performance of a matcher is evaluated for one group, matching at least one of the tuples in a pair. In pairwise fairness though, a pair of groups is considered for evaluation. Given a group of interest, the logic layer summarizes the workload into a confusion matrix, which is used in computing various fairness measures. Next, proper fairness measures are used to evaluate the EM task depending on the context of the task, the fairness paradigm of interest, etc. FairEM360 offers a set of 5 group fairness definitions specific to EM that users can use to evaluate their tasks. Finally, each group’s unfairness is calculated using subtraction-based or division-based notations of disparity and if the unfairness exceeds the fairness threshold specified by the user, the matcher is considered unfair for that group. For more details on fairness definitions, and unfairness measurement please refer to  (Shahbazi et al., 2023a).
@@ 2.3. Presentation Layer
The presentation layer determines whether a matcher w.r.t. a measure/group. To allow users to explore potential explanations for unfairness, FairEM360 provides three perspectives:
@@
  • Group-representation based Explanation. Responsible training in EM techniques requires access to unbiased data with proper representation of different groups and possible cases (Shahbazi et al., 2023b). over/under representation of different groups can bias the models in favor of some of the groups, making the model unfair. In particular, given the class imbalance nature of EM tasks, it is important to ensure proper representatives from different groups in both (match/unmatch) classes. This group of explanations analyzes the representation of each group within the dataset and conditioned on the matching ground truths.
@@ 2.4. Implementation Details
Figure 2. data import step.
Figure 3. matcher selection step.
Figure 6. ensemble-based resolutions.
@@ 3. Demonstration Scenarios
By clicking on each bar, FairEM360 provides insights into various aspects of the group, including its representation (coverage) in the training data, a random sample of problematic pairs associated with the measure and group, investigating the subgroup hierarchy unfairness if applicable (specific to setwise or intersectional sensitive attributes), group’s confusion matrix and factors that have contributed to the unfairness, among other details. Figure 7 provides an example of the potential explanations for a matcher’s unfair behavior towards a group. In this case the matcher has been unfair towards cn group w.r.t. True Positive Rate Parity. By examining the example-based explanations, one can infer that the matcher is erroneously matching entities from the cn group due to inherent similarities present in Chinese names compared to German names.

### v2

@@ Abstract.
Entity matching is one of the earliest tasks that occur in the big data pipeline and is alarmingly exposed to unintentional biases that affect the quality of data. Identifying and mitigating the biases that exist in the data or are introduced by the matcher at this stage can contribute to promoting fairness in downstream tasks. This demonstration showcases FairEM360, a framework for 1) auditing the output of entity matchers across a wide range of fairness measures and paradigms, 2) providing potential explanations for the underlying reasons for unfairness, and 3) providing resolutions for the unfairness issues through an exploratory process with human-in-the-loop feedback, utilizing an ensemble of matchers. We aspire for FairEM360 to contribute to the prioritization of fairness as a key consideration in the evaluation of EM pipelines.
Authors contributed equally to this work.
@@ 1. Introduction
FairEM360 incorporates a comprehensive set of group fairness definitions tailored for single and pairwise fairness metrics specifically designed for EM audits (Shahbazi et al., 2023a). FairEM360 also features a ready-to-use ensemble of 10 matchers for performing matching tasks. Furthermore, FairEM360 offers a diverse range of subgroup-based, measure-based, and group-representation based explanations that shed light on the reasoning behind matcher’s unfair behavior. Finally, through an exploratory process with human-in-the-loop feedback that navigates the combinatorial space of matchers, and fairness measures, FairEM360 is directed toward satisfying the fairness constraints required by the user while satisfying a minimum matching performance for various groups. To the best of our knowledge, FairEM360 stands out as the first system for responsible EM. The closest existing system to ours is FairER (Fanourakis et al., 2023), a framework for fair and explainable entity resolution. However, their definition of fairness differs from ours, and they only provide explanations on the matching outcomes, not on the unfairness. Similarly, systems such as LEMON (Barlaug, 2022), WYM (Baraldi et al., 2023), and ExplainER (Ebaid et al., 2019) focus on the explainability of the matching task. Fairness in EM has been studied in various settings: the impact of tuning matching thresholds to achieve fairness  (Moslemi and Milani, 2024), group-based training for different ethnicities to improve both accuracy and fairness in SVM-based ER (Makri et al., 2022), AUC-based fairness for EM/ER tasks and resolving the unfairness through a data augmentation (Nilforoushan et al., 2022), an experimental analysis of fairness in EM (Shahbazi et al., 2023a) and promoting fairness in the data preparation step in EM (Shahbazi et al., 2024).
@@ 2.1. Data Layer
Group Extraction: The first step in auditing an entity matcher for fairness is identifying meaningful groups/subgroups (e.g., white-male, black-female, etc.) from the sensitive attributes (e.g., race, sex), based on which the matcher should be audited. Sensitive attributes are attributes for which a matcher is likely to exhibit bias. Sensitive attributes are determined by a human expert. Depending on the type, cardinality, and number of sensitive attributes, FairEM360 navigates the space of all possible (sub)groups. It unifies the group representations using one-hot encoding, which enables representing the subgroups of individual entities and pairs of entities across various settings such as binary, non-binary, and setwise sensitive attributes. Entity encodings are the output of the data layer and will be passed as an input to the logic layer, where the fairness of a matcher is investigated for a group.
Preprocessing: If the user opts to utilize the integrated matchers in their evaluation, the data layer will be responsible for preprocessing the input datasets to ensure their compatibility with the matchers, as well as splitting the data into test, train, and validation sets. Our framework aligns with the format of established benchmark datasets, such as Magellan (Konda, 2018) and WDC (Primpeli et al., 2019). The datasets are passed to the logic layer to be trained.
@@ 2.2. Logic Layer
The actual evaluation of the output from an EM task occurs in the logic layer. The input to the logic layer is a workload of tuples each consisting of left and right entity groups (extracted in the data layer), the matcher prediction, and the matching ground truth. A workload is a test set of tuples for evaluating an entity matcher. Given the pairwise nature of EM tasks, single fairness or pairwise fairness paradigms can be used to audit matchers for fairness. In single fairness, the performance of a matcher is evaluated for one group, matching at least one of the tuples in a pair. In pairwise fairness though, a pair of groups is considered for evaluation. Given a group of interest, the logic layer summarizes the workload into a confusion matrix, which is used in computing various fairness measures. Next, proper fairness measures are used to evaluate the EM task depending on the context of the task, the fairness paradigm of interest, etc. FairEM360 offers a set of 5 group fairness definitions specific to EM that users can use to evaluate their tasks. Finally, each group’s unfairness is calculated using subtraction-based or division-based notations of disparity and if the unfairness exceeds the fairness threshold specified by the user, the matcher is considered unfair for that group. In the following, we provide more details on fairness definitions and unfairness measurement:
The input to the logic layer is a workload $w$ of $n$ tuples each having a correspondence $t=(e_{i},e_{j},h,y)$, where $h$ is a binary variable indicating the result of entity matching (match or non-match) for entities with encodings $e_{i}$ and $e_{j}$, and $y$ is a binary variable indicating the ground-truth for matching. A workload is defined as a test set of tuples for evaluating an entity matcher.
Fairness Paradigms: Given the pairwise nature of EM tasks, there are two paradigms to audit matchers for fairness. In the single fairness, the performance of a matcher is evaluated for one subgroup $s$ against either entity in a pair. Given a correspondence $(e_{i},e_{j},h,y)$ and a subgroup $s$ of interest, FairEM360 considers the correspondence legitimate, if either $e_{i}$ or $e_{j}$ belong to subgroups $s$. In pairwise fairness, the performance of a matcher is evaluated for a pair of subgroups $s,s^{\prime}$ against both entities in a pair. Given a correspondence $(e_{i},e_{j},h,y)$ and a pair of sub-groups $s,s^{\prime}$ of interest, FairEM360 considers the correspondence legitimate, if either $e_{i}$ or $e_{j}$ belong to subgroups $c$. if $e_{i}$ belongs to $s$ and $e_{j}$ belongs to $s^{\prime}$, or vice versa.
Given a subgroup of interest, the logic layer summarizes the workload $w$ into a confusion matrix, which is later used in computing various fairness measures. When auditing with single and pairwise fairness, each correspondence $(e_{i},e_{j},m,y)$ corresponds either to True Positives (TPs), False Positives (FPs), False Negatives (FNs), or True Negatives (TNs). Note that the result is counted both for the group(s) of $e_{i}$ and the group(s) of $e_{j}$. This is unlike regular classification where every row corresponds to one entity and hence is counted once.
Fairness Measures: Depending on the context of an EM task at hand, proper fairness measures should be employed. Besides, a major difference between EM and regular classification tasks is that the input to EM tasks is a pair of records. Due to its pairwise matching nature, class imbalance is a distinguishing property of EM, compared to regular classification tasks. In these settings, the fairness measure for successfully discovering these events is Positive Predictive Value Parity (PPVP). Another important measure in this context is True Positive Rate Parity (TPRP), a.k.a Equal Opportunity, which focuses on correct match predictions among the (rare) true matches. For further information regarding the fairness measures in the context of EM, refer to  (Shahbazi et al., 2023a).
Measuring Unfairness: Consider a fairness notion and a subgroup $g_{i}\in\mathbf{g}$. In a perfect situation, the matcher should satisfy the parity (equality) between two probabilities in the following form:
(1) $\forall g_{i}\in\mathbf{g},Pr(\alpha\leavevmode\nobreak\ |\leavevmode\nobreak\ \beta,g_{i})=Pr(\alpha\leavevmode\nobreak\ |\leavevmode\nobreak\ \beta)$
where $\alpha$ and $\beta$ are specified by the fairness measure. For example, for Positive Predictive Parity, $\alpha$ is $y=`M\textrm{'}$ ’ and $\beta$ is $h(e,e^{\prime})=`M\textrm{'}$ ’.
On the other hand, due to the trade-offs between different fairness notions and the impossibilities theorems, it is often not possible to satisfy complete parity on all fairness measures. As a result, considering a threshold value (e.g. the 20% rule suggests the threshold as $0.2$), the objective is to make sure that disparity (a.k.a unfairness) is less than the threshold. Given a fairness notion and a subgroup $g_{i}\in\mathbf{g}$, the disparity can be computed using subtraction, as follows:
(2) $F^{(s)}_{\alpha,\beta}(g_{i})=\max\Big{(}0\leavevmode\nobreak\ ,\leavevmode\nobreak\ Pr(\alpha\leavevmode\nobreak\ |\leavevmode\nobreak\ \beta)-Pr(\alpha\leavevmode\nobreak\ |\leavevmode\nobreak\ \beta,g_{i})\Big{)}$
Alternatively, given a fairness notion and a subgroup $g_{i}\in\mathbf{g}$, the disparity can be computed using division, as follows:
(3) $F^{(d)}_{\alpha,\beta}(g_{i})=\max\Big{(}0\leavevmode\nobreak\ ,1-\frac{Pr(\alpha\leavevmode\nobreak\ |\leavevmode\nobreak\ \beta,g_{i})}{Pr(\alpha\leavevmode\nobreak\ |\leavevmode\nobreak\ \beta)}\Big{)}$
In FairEM360, we provide the choice to utilize either subtraction-based and division-based notation of disparity.
@@ 2.3. Presentation Layer
Multiple-workload Analysis: We also consider the scenario where multiple instances of test data a.k.a workloads are available from a matcher, i.e. different test data may become available at different times or different samples from the underlying test data distribution may be available. In this case, the logic layer evaluates workloads separately. For $k$ workloads, the input to the presentation layer is a population of $k$ disparity values for each group and measure combination. A population for group $g$ and measure $m$ includes the disparity of $g$ in every workload w.r.t. $m$. To assess the fairness of a matcher on group $g$ using measure $m$ and $k$ workloads, we employ the standard hypothesis testing. For a group $g$ and measure $m$, the fairness hypothesis testing considers the null hypothesis that the matcher is fair on $g$ and the alternative hypothesis that the matcher is unfair on $g$. For more than one workload, FairEM360 chooses the appropriate $z$-test statistics. Next, the test statistics and corresponding $p$-value are computed as the probability of getting the observed test statistic or something more extreme when the null hypothesis is true. Finally, given a significance level $\alpha$, the null hypothesis in favor of the alternative is rejected if $\alpha\leq p$-value, or not if $\alpha>p$-value. If a user chooses to perform multiple-workload analysis on a single provided test data set, FairEM360 generates $k$ workloads by random sampling with replacement from the data set. The goal of this analysis would be to verify whether an unfair group happened by chance or is indeed repeatable.
@@
The presentation layer determines whether a matcher w.r.t. a measure/group. To allow users to explore potential explanations for unfairness, FairEM360 provides four perspectives:
@@
  • Group-representation based Explanation. Responsible training in EM techniques requires access to unbiased data with proper representation of different groups and possible cases (Shahbazi et al., 2023b). Over/under-representation of different groups can bias the models in favor of some of the groups, making the model unfair. In particular, given the class imbalance nature of EM tasks, it is important to ensure proper representatives from different groups in both (match/unmatch) classes. This group of explanations analyzes the representation of each group within the dataset and conditioned on the matching ground truths.
@@ 3. Demonstration Scenarios
By clicking on each bar, FairEM360 provides insights into various aspects of the group, including its representation (coverage) in the training data, a random sample of problematic pairs associated with the measure and group, investigating the subgroup hierarchy unfairness if applicable (specific to setwise or intersectional sensitive attributes), group’s confusion matrix and factors that have contributed to the unfairness, among other details. Figure 7 provides an example of the potential explanations for a matcher’s unfair behavior towards a group. In this case the matcher has been unfair towards cn group w.r.t. True Positive Rate Parity. By examining the example-based explanations, one can infer that the matcher is erroneously matching entities from the cn group due to inherent similarities present in Chinese names compared to German names. Note that any dataset with any grouping of data for which we require equal performance of the matcher can be evaluated by FairEM360. It does not necessarily need to be social data; however, in the demo, we focus only on such datasets. For analysis on non-social benchmark datasets, please refer to (Shahbazi et al., 2023a).
@@ 4. Conclusion
Acknowledgement
This work was supported in part by NSF grants 2107290 and 2348919.